OpenStudy (anonymous):

on a factory floor, 40 out of every 150 toy robots is defective. What percent of the toy robots are defective? Round your answer to the nearest hundreth:) Answer Choices: A: 60% B: 2.67% C: 0.6% D:26.67%

OpenStudy (gudden):

percentage of those defective = (no of defective / total)*100 = (40/150 )100 = 26.67%

OpenStudy (anonymous):

Thanx so much ! i aprectiate it :)